OpenStudy (tomfoolery1):

Im struggling! WILL MEDAL Use the given degree of confidence and sample data to construct a confidence interval for the population mean µ. Assume that the population has a normal distribution. n = 10, x = 12.3, s = 3.9, 95% confidence 9.55 < µ < 15.05 9.52 < µ < 15.08 9.51 < µ < 15.09 10.04 < µ < 14.56

OpenStudy (math&ing001):

What does 's' represent ?

OpenStudy (tomfoolery1):

i believe that it must mean ST. DEV

OpenStudy (math&ing001):

Ok, then ! Give me a few moments I'm a little rusty on this =)

OpenStudy (tomfoolery1):

Thank you! I am so grateful! =)

OpenStudy (math&ing001):

95% confidence means that 95% of values are within 2 standard deviations of the mean (see attached) A confidence interval for \(\mu\) is given by : \[x \pm 2\sigma _{x}\]\[12.3 \pm 2 \sqrt{\frac{ 3.9 }{ 10 }}\] Reference: http://www.kean.edu/~fosborne/bstat/06amean.html

OpenStudy (tomfoolery1):

11.051 - 13.549 is what i got. do i subtract 2 form either side?

OpenStudy (math&ing001):

You got that how ?

OpenStudy (math&ing001):

\[12.3 \pm 2 \sqrt{\frac{ 3.9 }{ 10 }}\] Translates to : \[12.3 - 2 \sqrt{\frac{ 3.9 }{ 10 }} <\mu<12.3 + 2 \sqrt{\frac{ 3.9 }{ 10 }}\]

OpenStudy (tomfoolery1):

by computing the last calculation 12.3+2sqrt3.9/10

OpenStudy (math&ing001):

Little mistake there, since s is the standard deviation, there's no need for the square root on \(\sigma\). Hence \[12.3 - 2 \frac{ 3.9 }{\sqrt{ 10 }} <\mu<12.3 + 2 \frac{ 3.9 }{\sqrt{ 10 }}\]